# Effects of intertube dipole-dipole interactions in nearly integrable one-dimensional 162Dy gases

### v1

@@ V Other initial states
Next, we explore how the modeling with and without the intertube DDI correction compare to one another for other initial states experimentally created in Ref. [42]. We consider states with $\theta_{B}=55^{\circ}$ ($g_{\rm 1D}^{\rm DDI}=0$) but different values of $g_{\rm 1D}^{\rm vdW}$ in Sec. V.1, and then states with different $\theta_{B}$ in Sec. V.2.
@@ V.1 Fixed $\theta_{B}=55^{\circ}$ and different $g_{\rm 1D}^{\rm vdW}$
We first consider the case in which after the experimental sequence studied in Sec. III is completed, the magnitude of the magnetic field is changed. As a result, the contact interaction $g^{\rm vdW}_{\rm 1D}$ changes via a Feshbach resonance. We consider the two experimental cases studied in Ref. [42], with $g^{\rm vdW}_{\rm 1D}=4.1$ and $20.4$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ resulting in a final effective $\gamma_{T}=3.1$ (3.1) and 17 (16), respectively, in the presence (absence) of the intertube DDI. Since $\theta_{B}$ is kept fixed at $\theta_{B}=55^{\circ}$, the intratube DDI remains near zero and the intertube DDI is not different from our previous example.
@@
Figure 7: Temperature distribution for different values of $g_{\rm 1D}^{\rm vdW}$. $N_{T}/N_{\text{total}}$ for other initial states with $\theta_{B}=55^{\circ}$ and: (a) $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ and (b) $g^{\rm vdW}_{\rm 1D}=20.4$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$. Experimentally, $g^{\rm vdW}_{\rm 1D}$ is changed after loading into the deep 2D lattice by adiabatically modifying the magnitude of the magnetic field $B$ while keeping fixed its direction $\theta_{B}=55^{\circ}$. In our modeling, we modify $T$ in intervals of $0.5$ nK to find the closest entropy in each tube to that at decoupling.
In Fig. 7, we show our results for the distribution of temperatures of the states with $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ [Fig. 7(a)] and $g^{\rm vdW}_{\rm 1D}=20.4$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ [Fig. 7(b)]. Changing $g^{\rm vdW}_{\rm 1D}$ in the range of values experimentally explored does not significantly change the temperature of the tubes. The largest changes in the temperatures are due to the dimensional reduction from three dimensions to one dimension, and due to lowering $f_{x}$ at the end of the loading process. We find that reducing $g^{\rm vdW}_{\rm 1D}$ slightly lowers the temperature of the tubes. As for the case discussed in Sec. III.2 in which $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$, we also find that the intertube DDI slightly lowers the temperature, and that its effect is more pronounced for the tubes with a lower number of atoms.
@@
Figure 8: Density and rapidity distributions for different values of $g_{\rm 1D}^{\rm vdW}$. (a),(c) Density and (b),(d) rapidity distributions for (a),(b) $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ and (c),(d) $g^{\rm vdW}_{\rm 1D}=20.4$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$. The curves plotted for each distribution are the same as in Fig. 6.
@@
In Fig. 8, we show the density $n(x)$ [Figs. 8(a) and 8(c)] and rapidity $f(\theta)$ [Figs. 8(b) and 8(d)] distributions for the states with $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ [Figs. 8(a) and 8(b)] and $g^{\rm vdW}_{\rm 1D}=20.4$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ [Figs. 8(c) and 8(d)]. As expected, reducing $g^{\rm vdW}_{\rm 1D}$ results in a narrowing of both the density and the rapidity distributions. The results are qualitatively similar to those for $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ in Sec. III.2. With the intertube DDI correction, the density and rapidity distributions in the initial state are slightly narrower than in the absence of the correction and, after the expansion, the rapidity distribution with the DDI correction is nearly indistinguishable from that in the absence of the correction. As discussed in Ref. [42], the rapidity distributions from our simulations approach the ones measured experimentally as $g^{\rm vdW}_{\rm 1D}$ ($\gamma_{T}$) increases.
@@ V.2 Fixed $g_{\rm 1D}^{\rm vdW}$ and different $\theta_{B}$
Finally, we consider the case in which after the experimental sequence studied in Sec. III is completed, the magnetic field is rotated without change in amplitude. The contact interaction $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ does not change with the rotation of the magnetic field, but both the intratube (which becomes nonzero) and the intertube DDI change. To account for the intratube part of DDI, we use the approximation in Eq. (II.2), i.e., we treat it as a correction $g^{\rm DDI}_{\rm 1D}$ to the contact interaction. Three states with different $\theta_{B}\neq 55^{\circ}$ were prepared in Ref. [42] using that sequence: (i) one with $\theta_{B}=90^{\circ}$ resulting in an additional effective repulsive contact interaction with $g^{\rm DDI}_{\rm 1D}=2.7$, (ii) one with $\theta_{B}=35^{\circ}$ resulting in an additional effective attractive contact interaction with $g^{\rm DDI}_{\rm 1D}=-2.7$, and (iii) one with $\theta_{B}=0^{\circ}$ resulting in an additional effective attractive contact interaction with $g^{\rm DDI}_{\rm 1D}=-5.4$. The corresponding values of $\gamma_{T}$ in the presence (absence) of the intertube DDI are found to be: (i) $\gamma_{T}=2.2$ (2.1), (ii) $\gamma_{T}=4.5$ (4.5), and (iii) $\gamma_{T}=8.6$ (8.5).
Figure 9: Temperature distribution for different values of $\theta_{B}$. $N_{T}/N_{\text{total}}$ for initial states with $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ and: (a) $\theta_{B}=0^{\circ}$, (b) $\theta_{B}=35^{\circ}$, and (c) $\theta_{B}=90^{\circ}$. Experimentally, $\theta_{B}$ is changed adiabatically (keeping the magnitude of magnetic field $B$ constant) after the loading in the deep 2D lattice. In our modeling, we modify $T$ in intervals of $0.5$ nK to find the closest entropy in each tube to that at decoupling.
@@
In Fig. 9, we show our results for the distribution of temperatures after the rotation of the magnetic field. They are qualitatively similar to those reported for $\theta_{B}=55^{\circ}$. In the presence of the intertube DDI correction, the temperatures are slightly lower and more widely distributed than in its absence. Changing $\theta_{B}$ slightly shifts those distributions resulting in an average temperature that increases slightly with increasing $\theta_{B}$.
@@
Figure 10: Density and rapidity distributions for different values of $\theta_{B}$. (a),(c),(e) Density and (b),(d),(f) rapidity distributions for (a),(b) $\theta_{B}=0^{\circ}$, (c),(d) $\theta_{B}=35^{\circ}$, and (e),(f) $\theta_{B}=90^{\circ}$. The curves plotted for each distribution are the same as in Figs. 6 and 8.
@@
In Figs. 10(a), 10(c), and 10(e) we plot the density distribution without and with the intertube DDI correction for $\theta_{B}=0^{\circ}$, $35^{\circ}$, and $90^{\circ}$, respectively. The effect of the intertube DDI can be seen to be small in all cases. As we mentioned earlier, the intertube DDI produces two competing effects on the density distribution, it tends to broaden the density distribution due to antitrapping, while it tends to narrow the density distribution by lowering the temperature of the tubes. The competition between those two effects results in a density distribution in the presence of the intertube DDI that is slightly broader for $\theta_{B}=0^{\circ}$ [Fig. 10(a)], nearly equal for $\theta_{B}=35^{\circ}$ [Fig. 10(c)], and slightly narrower for $\theta_{B}=90^{\circ}$ [Fig. 10(e)] than those in the absence of the intertube DDI.
@@
In Figs. 10(b), 10(d), and 10(f), we plot the rapidity distributions with and without the intertube DDI correction for $\theta_{B}=0^{\circ}$, $35^{\circ}$, and $90^{\circ}$, respectively. The intertube DDI can be seen to always result in a narrowing of the rapidity distribution with respect to that in the absence of the DDI, and this “narrowing” effect becomes more prominent as $\theta_{B}$ decreases. During the expansion, all the rapidity distributions broaden and, again, they become very similar to those in the absence of the intertube DDI. Overall, decreasing $\theta_{B}$ results in narrower rapidity distributions both in the modeling and in the experimentally measured rapidity distribution. This feature is properly captured by our modeling, though the peak heights continue to not be accurately simulated. In addition, we note that the agreement between the results of our modeling and the experimental results improve as $\theta_{B}$ decreases. There is no a priori reason for this to be the case, so we expect the observed improvement in the agreement to be the result of a cancellation of effects that were not accounted for in our modeling.
VI Summary
@@
We systematically studied the effects of the intertube DDI in recent experiments with arrays of nearly integrable 1D dipolar Bose gases of 162Dy atoms. The leading-order effect of the intertube DDI is accounted for by treating it as a self-consistent modification to the 1D confining potentials, both during the state preparation and during the rapidity measurements. Our results suggest that this correction changes only marginally the prediction for the measured rapidity distribution. The rapidity distributions obtained by adding the intertube DDI correction are very close to those obtained in its absence because of the nearly perfect cancellation of the effect of the intertube DDI in the initial state with that of the effect of it during the expansion. In the former, it leads to a narrowing of the distributions, while in the latter it leads to a broadening, and those two effects are coincidentally nearly identical in our setup. We also found that the intertube DDI slightly cools the 1D gases by causing an effective antitapping effect. The cooling is more pronounced in the tubes with lower atomic filling lying away from the center of the 2D array of 1D gases.
@@
Having found that the disagreements between the simulations and the experimental results in Ref. [42] are unlikely to be related to the intertube DDI, we strongly suspect that they are the result of nonthermal effects (due to the near-integrability of the 1D gases) after the tubes decouple. In future work, we plan to explore how to better model the decoupling process near integrability, as well as how to model the final stage of the loading process without the assumption that each 1D gas is in thermal equilibrium at a fixed entropy.
@@ Acknowledgements.
Y.Z. acknowledges support from the Dodge Family Postdoc Fellowship at the University of Oklahoma, M.R. acknowledges support from the NSF (PHY2309146), and B.L.L. acknowledges support from the NSF (PHY2308540) and AFOSR (FA9550-22-1-0366).
Appendix A Rounding the number of particles and the effect of tubes with very few particles
In the modeling in Ref. [42], we rounded the number of particles in each tube to the closest integer. This simplification was necessary because the quantum Monte Carlo simulations of the momentum distribution are computationally demanding and, therefore, it was not possible to simulate thousands of tubes with slightly different fillings. Because in this work we focus on the density and rapidity distributions, which are obtained using the TBA and the LDA, we can carry out the calculations without the rounding. In Fig. 11, we show the rapidity distributions obtained in Ref. [42] and now without rounding (both ignoring the intertube DDI) for the state after fully loading the atoms into the 2D lattice, as described in Sec. III. One can see that, because of the rounding, the rapidity distribution from Ref. [42] is slightly narrower and further away from the experimentally measured distribution. In consequence, in this work we carry out all our calculations without rounding the number of atoms in the tubes.
Figure 11: Rapidity distributions from this work and from Ref. [42]. Rapidity distributions for $\theta_{B}=55^{\circ}$ and $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$ at the end of state preparation in the absence of the intertube DDI. We show results for the average over all tubes (blue solid line) and over tubes with $N_{l}\geq 2$ (red dashed line) obtained in this work, and from the modeling used in Ref. [42] in which the number of atoms in each tube was rounded to the closest integer (green dashed-dotted line). The experimental rapidity distribution from Ref. [42] is shown as an orange solid line.
On the other hand, because of the self-consistent nature of the calculations carried out in this work to account for the intertube DDI, it is very challenging to include the thousands of tubes that have less than two atoms. As mentioned in the main text, those account for less than 5% of the atoms. In Fig. 11 we show the normalized rapidity distribution obtained after ignoring those tubes. It is very close to that in which those tubes are included (rounding to the closest integer has a stronger effect). Since our main goal is to contrast the results obtained without and with the intertube DDI, we remove those tubes from the calculations (without and with intertube DDI) used to obtain the results reported in Sec. III.2 and in later sections.

### v2

@@ V Other initial states
Next we explore how the modeling with and without the intertube DDI correction compare to one another for other initial states experimentally created in Ref. [43]. We consider states with $\theta_{B}=55^{\circ}$ ($g_{\rm 1D}^{\rm DDI}=0$) but different values of $g_{\rm 1D}^{\rm vdW}$ in Sec. V.1, and then states with different $\theta_{B}$ in Sec. V.2.
@@ V.1 Fixed $\theta_{B}=55^{\circ}$ and different $g_{\rm 1D}^{\rm vdW}$
We first consider the case in which after the experimental sequence studied in Sec. III is completed, the magnitude of the magnetic field is changed. As a result, the contact interaction $g^{\rm vdW}_{\rm 1D}$ changes via a Feshbach resonance. We consider the two experimental cases studied in Ref. [43], with $g^{\rm vdW}_{\rm 1D}=4.1$ and $20.4$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ resulting in a final effective $\gamma_{T}=3.1$ (3.1) and 17 (16), respectively, in the presence (absence) of the intertube DDI. Since $\theta_{B}$ is kept fixed at $\theta_{B}=55^{\circ}$, the intratube DDI remains near zero and the intertube DDI is not different from our previous example.
@@
Figure 7: Temperature distribution for different values of $g_{\rm 1D}^{\rm vdW}$. Here $N_{T}/N_{\rm tot}$ is plotted for other initial states with $\theta_{B}=55^{\circ}$ and (a) $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ and (b) $g^{\rm vdW}_{\rm 1D}=20.4$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$. Experimentally, $g^{\rm vdW}_{\rm 1D}$ is changed after loading into the deep 2D lattice by adiabatically modifying the magnitude of the magnetic field $B$ while keeping fixed its direction $\theta_{B}=55^{\circ}$. In our modeling, we modify $T$ in intervals of $0.5$ nK to find the closest entropy in each tube to that at decoupling.
@@
Figure 8: Density and rapidity distributions for different values of $g_{\rm 1D}^{\rm vdW}$. (a) and (c) Density and (b) and (d) rapidity distributions for (a) and (b) $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ and (c) and (d) $20.4$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$.
In Fig. 7 we show our results for the distribution of temperatures of the states with $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ [Fig. 7(a)] and $g^{\rm vdW}_{\rm 1D}=20.4$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ [Fig. 7(b)]. Changing $g^{\rm vdW}_{\rm 1D}$ in the range of values experimentally explored does not significantly change the temperature of the tubes. The largest changes in the temperatures are due to the dimensional reduction from three dimensions to one dimension and due to lowering $f_{x}$ at the end of the loading process. We find that reducing $g^{\rm vdW}_{\rm 1D}$ slightly lowers the temperature of the tubes. As for the case discussed in Sec. III.2 in which $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$, we also find that the intertube DDI slightly lowers the temperature and that its effect is more pronounced for the tubes with a lower number of atoms.
@@
In Fig. 8 we show the density $n(x)$ [Figs. 8(a) and 8(c)] and rapidity $f(\theta)$ [Figs. 8(b) and 8(d)] distributions for the states with $g^{\rm vdW}_{\rm 1D}=4.1$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ [Figs. 8(a) and 8(b)] and $g^{\rm vdW}_{\rm 1D}=20.4$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ [Figs. 8(c) and 8(d)]. As expected, reducing $g^{\rm vdW}_{\rm 1D}$ results in a narrowing of both the density and the rapidity distributions. The results are qualitatively similar to those for $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ in Sec. III.2. With the intertube DDI correction, the density and rapidity distributions in the initial state are slightly narrower than in the absence of the correction and, after the expansion, the rapidity distribution with the DDI correction is nearly indistinguishable from that in the absence of the correction. As discussed in Ref. [43], the rapidity distributions from our simulations approach the ones measured experimentally as $g^{\rm vdW}_{\rm 1D}$ ($\gamma_{T}$) increases.
@@ V.2 Fixed $g_{\rm 1D}^{\rm vdW}$ and different $\theta_{B}$
Finally, we consider the case in which after the experimental sequence studied in Sec. III is completed, the magnetic field is rotated without change in amplitude. The contact interaction $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ does not change with the rotation of the magnetic field, but both the intratube (which becomes nonzero) and the intertube DDI change. To account for the intratube DDI, we use the approximation in Eq. (II.2), i.e., we treat it as a correction $g^{\rm DDI}_{\rm 1D}$ to the contact interaction. Three states with different $\theta_{B}\neq 55^{\circ}$ were prepared in Ref. [43] using that sequence: (i) one with $\theta_{B}=90^{\circ}$ resulting in an additional effective repulsive contact interaction with $g^{\rm DDI}_{\rm 1D}=2.7$, (ii) one with $\theta_{B}=35^{\circ}$ resulting in an additional effective attractive contact interaction with $g^{\rm DDI}_{\rm 1D}=-2.7$, and (iii) one with $\theta_{B}=0^{\circ}$ resulting in an additional effective attractive contact interaction with $g^{\rm DDI}_{\rm 1D}=-5.4$. The corresponding values of $\gamma_{T}$ in the presence (absence) of the intertube DDI are (i) $\gamma_{T}=2.2$ (2.1), (ii) $4.5$ (4.5), and (iii) $8.6$ (8.5).
@@
In Fig. 9 we show our results for the distribution of temperatures after the rotation of the magnetic field. They are qualitatively similar to those reported for $\theta_{B}=55^{\circ}$. In the presence of the intertube DDI correction, the temperatures are slightly lower and more widely distributed than in its absence. Changing $\theta_{B}$ slightly shifts those distributions resulting in an average temperature that increases slightly with increasing $\theta_{B}$.
Figure 9: Temperature distribution for different values of $\theta_{B}$. Here $N_{T}/N_{\rm tot}$ is plotted for initial states with $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ and (a) $\theta_{B}=0^{\circ}$, (b) $\theta_{B}=35^{\circ}$, and (c) $\theta_{B}=90^{\circ}$. Experimentally, $\theta_{B}$ is changed adiabatically (keeping the magnitude of magnetic field $B$ constant) after loading into the deep 2D lattice. In our modeling, we modify $T$ in intervals of $0.5$ nK to find the closest entropy in each tube to that at decoupling.
@@
Figure 10: Density and rapidity distributions for different values of $\theta_{B}$. (a), (c), and (e) Density and (b), (d), and (f) rapidity distributions for (a) and (b) $\theta_{B}=0^{\circ}$, (c) and (d) $35^{\circ}$, and (e) and (f) $90^{\circ}$.
@@
In Figs. 10(a), 10(c), and 10(e) we plot the density distribution without and with the intertube DDI correction for $\theta_{B}=0^{\circ}$, $35^{\circ}$, and $90^{\circ}$, respectively. The effect of the intertube DDI is small in all cases. As we mentioned earlier, the intertube DDI produces two competing effects on the density distribution: It tends to broaden the density distribution due to antitrapping, while it tends to narrow the density distribution by lowering the temperature of the tubes. The competition between those two effects results in a density distribution in the presence of the intertube DDI that is slightly broader for $\theta_{B}=0^{\circ}$ [Fig. 10(a)], nearly equal for $\theta_{B}=35^{\circ}$ [Fig. 10(c)], and slightly narrower for $\theta_{B}=90^{\circ}$ [Fig. 10(e)] than those in the absence of the intertube DDI.
@@
In Figs. 10(b), 10(d), and 10(f) we plot the rapidity distributions with and without the intertube DDI correction for $\theta_{B}=0^{\circ}$, $35^{\circ}$, and $90^{\circ}$, respectively. The intertube DDI can be seen to always result in a narrowing of the rapidity distribution with respect to that in the absence of the DDI, and this narrowing effect becomes more prominent as $\theta_{B}$ decreases. During the expansion, all the rapidity distributions broaden and again they become very similar to those in the absence of the intertube DDI. Overall, decreasing $\theta_{B}$ results in narrower rapidity distributions both in the modeling and in the experimentally measured rapidity distribution. This feature is properly captured by our modeling, though the peak heights continue to not be accurately simulated. In addition, we note that the agreement between the results of our modeling and the experimental results improve as $\theta_{B}$ decreases. There is no a priori reason for this to be the case, so we expect the observed improvement in the agreement to be the result of a cancellation of effects that were not accounted for in our modeling.
VI Summary and outlook
@@
We systematically studied the effects of the intertube DDI in recent experiments with arrays of nearly integrable 1D dipolar Bose gases of 162Dy atoms. The leading-order effect of the intertube DDI is accounted for by treating it as a self-consistent modification to the 1D confining potentials, both during the state preparation and during the rapidity measurements. Our results suggest that this correction changes only marginally the prediction for the measured rapidity distribution. The rapidity distributions obtained by adding the intertube DDI correction are very close to those obtained in its absence because of the nearly perfect cancellation of the effect of the intertube DDI in the initial state with that during the expansion. In the former it leads to a narrowing of the distributions, while in the latter it leads to a broadening, and those two effects are coincidentally nearly identical in our setup. We also found that the intertube DDI slightly cools the 1D gases by generating an effective antitrap. The cooling is more pronounced in the tubes with lower atomic filling lying away from the center of the 2D array of 1D gases.
@@
Having found that the disagreements between the simulations and the experimental results in Ref. [43] are unlikely to be related to the intertube DDI, we strongly suspect that they are the result of nonthermal effects (due to the near-integrability of the 1D gases) after the tubes decouple. In future work, we plan to explore how to better model the decoupling of the tubes into independent 1D gases. There is no reason for all the tubes to decouple at the same lattice depth $U^{*}_{\rm 2D}$, and different decoupling depths depending on the filling and location of the tubes may result in a distribution of temperatures at decoupling. Since we found our modeling to be sensitive to the temperature $T^{*}_{\rm 2D}$ at decoupling, accounting for a distribution of temperatures $T^{*}_{\rm 2D}$ should provide a better description of the experiments. Also, after decoupling, the 1D gases are nearly integrable, so we will explore descriptions of the ensuing adiabatic processes that take place in the experiments without assuming that each 1D gas is in thermal equilibrium throughout those processes.
@@ Acknowledgements.
Y.Z. acknowledges support from the Dodge Family Postdoc Fellowship at the University of Oklahoma, M.R. acknowledges support from the NSF (Grant No. PHY2309146) and B.L.L. acknowledges support from the NSF (Grant No. PHY2308540) and AFOSR (Grant No. FA9550-22-1-0366).
Appendix: Rounding the number of particles and the effect of tubes with very few particles
In the modeling in Ref. [43], we rounded the number of particles in each tube to the closest integer. This simplification was necessary because the quantum Monte Carlo simulations of the momentum distribution are computationally demanding and therefore it was not possible to simulate thousands of tubes with slightly different fillings. Because in this work we focus on the density and rapidity distributions, which are obtained using the TBA and the LDA, we can carry out the calculations without the rounding. In Fig. 11 we show the rapidity distributions obtained in Ref. [43] and now without rounding (both ignoring the intertube DDI) for the state after fully loading the atoms into the 2D lattice, as described in Sec. III. One can see that, because of the rounding, the rapidity distribution from Ref. [43] is slightly narrower and further away from the experimentally measured distribution. In consequence, in this work we carry out all our calculations without rounding the number of atoms in the tubes.
On the other hand, because of the self-consistent nature of the calculations carried out in this work to account for the intertube DDI, it is very challenging to include the thousands of tubes that have less than two atoms. As mentioned in the main text, those account for less than 5% of the atoms. In Fig. 11 we show the normalized rapidity distribution obtained after ignoring those tubes. It is very close to that in which those tubes are included (rounding to the closest integer has a stronger effect). Since our main goal is to contrast the results obtained without and with the intertube DDI, we remove those tubes from the calculations (without and with intertube DDI) used to obtain the results reported in Sec. III.2 and in following sections.
Figure 11: Rapidity distributions from this work and from Ref. [43]. Rapidity distributions for $\theta_{B}=55^{\circ}$ and $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$ at the end of state preparation in the absence of the intertube DDI. We show results for the average over all tubes (blue solid line) and over tubes with $N_{\ell}\geq 2$ (red dashed line) obtained in this work, and from the modeling used in Ref. [43] in which the number of atoms in each tube was rounded to the closest integer (green dash-dotted line). The experimental rapidity distribution from Ref. [43] is shown as an orange solid line.